# The Role of Learning Algorithms in Collective Action

### v2

@@ Abstract
Collective action in machine learning is the study of the control that a coordinated group can have over machine learning algorithms. While previous research has concentrated on assessing the impact of collectives against Bayes (sub)-optimal classifiers, this perspective is limited in that it does not account for the choice of learning algorithm. Classifiers seldom behave like Bayes classifiers and are influenced by the choice of learning algorithms along with their inherent biases. In this work, we initiate the study of how the choice of the learning algorithm plays a role in the success of a collective in practical settings. Specifically, we focus on distributionally robust optimization (DRO), popular for improving a worst group error, and on the ubiquitous stochastic gradient descent (SGD), due to its inductive bias for “simpler” functions. Our empirical results, supported by a theoretical foundation, show that the effective size and success of the collective are highly dependent on properties of the learning algorithm. This highlights the necessity of taking the learning algorithm into account when studying the impact of collective action in machine learning.
@@ 1 Introduction
With the rapid increase in deployed machine learning models, a large number of firms rely on data contributed by users (Gerlitz and Helmond, 2013) to train their algorithms. In response to this, consumers and users have searched for ways to alter their data to influence the outputs of such models (Chen, 2018; Burrell et al., 2019; Rahman, 2021). Algorithmic collective action (Olson, 1965; Hardt et al., 2023) has emerged as a formal framework to study the effect a coordinated group of individuals can have on such models, by altering the data they provide to the firm. This naturally leads to various technical questions regarding how large the collective needs to be, how effective different data altering strategies are, and which details about firm’s algorithm can be leveraged by the collective to achieve the best results.
@@
Hardt et al. (2023) initiated the formal study into the success of different collective action strategies. Their work provided theoretical analysis based on the fractional size of the collective and on the properties of the signal with regards to the original data distribution. However, their results only hold for Bayes (sub)-optimal classifiers which do not immediately adapt to peculiarities of practical learning algorithms. In this work, we investigate two types of commonly used learning algorithms that exhibit distinct properties: (1) Distributional robustness and (2) Simplicity bias. We show how these properties lead to significantly different levels of collective success that are unexplained by prior work.
@@
When the data is composed of multiple sub-populations, algorithms that optimise for average performance often perform poorly in minority sub-populations (Meinshausen and Bühlmann, 2015). A fairness-focused firm will want to ensure that their trained learning model performs well uniformly on all sub-populations, as opposed to being on-average good. Distributionally Robust Optimisation (DRO) is a family of algorithms designed to maximize this per-group accuracy (Hashimoto et al., 2018; Wang et al., 2020). We show that as a consequence, a small collective achieves higher success when the training algorithm performs DRO, compared to standard empirical risk minimization (ERM). Conversely, a large collective in the same settings achieves lower success, contradicting the expectation set by previous work regarding the effectiveness of large collectives.
@@
Second, most machine learning algorithms used today are based on some form of gradient descent (GD). Such algorithms, including the popular Stochastic Gradient Descent (SGD), Adam (Kingma and Ba, 2015), and RMSProp, exhibit a preference for learning functions that are “simpler". This inductive bias of GD algorithms is popularly referred to as simplicity bias (Kalimeris et al., 2019; Shah et al., 2020). In practice, this preference results in the model “overlooking” certain complex features. We demonstrate that these overlooked features can be leveraged by a collective to design a strategy that will gain higher success compared to what is possible on a Bayes optimal classifier.
@@ 2.1 Collective Action
While they are sufficient to characterise the success of the collective for a Bayes (sub)-optimal learner, various practically deployed algorithms show different behaviours, as is discussed in Section 3 and 4.
@@ 2.2 Distributionally Robust Optimisation
In Section 3, we inspect collective action on a set of learning algorithms that target Distributionally-Robust objectives (Delage and Ye, 2010; Sagawa et al., 2020; Duchi and Namkoong, 2021). Intuitively, these algorithms aim to learn classifiers that perform equally well on a set of distributions as opposed to any single one. Formally, they minimise the following objective
@@
where ${\mathcal{Q}}_{p}$ is an uncertainty set of distributions close to $p$ over which we want to control the risk, $\ell$ is the loss function, and $g_{\theta}$ is a function with parameters $\theta$. There are many possible definitions of the uncertainty set that the algorithm is controlling for. One possible choice is an $f$-divergence111 Definition 2 in the appendix. ball (Ali and Silvey, 1966; Csiszár, 1967).
@@
Thus, DRO algorithms are often employed when there is concern about performance on “similar” distributions or on subgroups of the data (Namkoong and Duchi, 2016; Duchi and Namkoong, 2019). In Section 3, we investigate how the success of the collective changes when minimising ${\mathcal{R}}_{\mathrm{dro}}(\theta)$ as opposed to performing simple Empirical Risk Minimisation (ERM).
@@ 5 Conclusion and Future Work
In this work we conducted a theoretical and empirical study on how the choice of learning algorithms affects the success of a collective action to plant a signal. We presented various approaches a collective can take to design their signal when they have knowledge about the learning algorithm. In particular, these approaches include maintaining a small collective size against algorithms like JTT and LfF, influencing the validation set against CVaR-DRO, and changing the complexity of the signal against SGD. While we have focused on these three algorithms, these phenomena are relevant for several popular algorithms. Common algorithms in the topic of domain adaptation (Koh et al., 2021), distribution shift (Hendrycks and Dietterich, 2019), and improving fairness (Berk et al., 2017) use some form of DRO algorithms. It would be interesting to consider the impact of other kinds of algorithms used to improve worst group performance including un-supervsied and self-supervised representation learning (Shi et al., 2023) and fairness-inducing in-processing (Prost et al., 2019) and post-processing algorithms (Ţifrea et al., 2024). In a similar vein, recent research has seen a surge in algorithms that aim to improve safety of outputs of generative models including techniques like adversarial training (Ziegler et al., 2022), DPO (Rafailov et al., 2024), and RLHF (Christiano et al., 2017). It is important to also consider the impact of these algorithms on the success of a well-designed collective.
@@
Algorithmic biases akin to simplicity bias are also exhibited by SGD training of various state-of-the art deep neural networks. This includes a texture bias for CNN (Hermann et al., 2020), an in-context bias for language models (Levine et al., 2022), and word-order biases in LSTMs and transformers (White and Cotterell, 2021). Our work discusses how these biases can affect the design of the collective signal and how successful they are expected to be. However, several algorithms also display a different kind of biases. For example, differentially private algorithms are unable to learn minority sub-populations (Bagdasaryan et al., 2019; Sanyal et al., 2022) and adversarially robust algorithms (Madry et al., 2018) are used to improve the smoothness of learned models. One direction for future work is to consider the impact of these algorithmic biases on the success of the collective.
@@
Finally, a third important direction of future research is understanding what level of information and influence is required by the collective to design their signal. For example, our results suggest that improving success against certain DRO algorithms requires less information about the learning algorithm compared to success against SGD style algorithms. However, for iterative re-weighting algorithms like CVar-DRO (Sagawa et al., 2020), it is important to influence both the validation set and the training set. Future work should investigate whether a uniformly randomly selected collective can be as powerful as a strategically chosen collective (e.g. in adversarial vulnerability (Paleka and Sanyal, 2023)). We hope that our work will inspire further research into practical and algorithm-dependent view on collective action.

### v3

@@ Abstract
Collective action in machine learning is the study of the control that a coordinated group can have over machine learning algorithms. While previous research has concentrated on assessing the impact of collectives against Bayes (sub-)optimal classifiers, this perspective is limited in that it does not account for the choice of learning algorithm. Since classifiers seldom behave like Bayes classifiers and are influenced by the choice of learning algorithms along with their inherent biases, in this work we initiate the study of how the choice of the learning algorithm plays a role in the success of a collective in practical settings. Specifically, we focus on distributionally robust optimization (DRO), popular for improving a worst group error, and on the ubiquitous stochastic gradient descent (SGD), due to its inductive bias for “simpler” functions. Our empirical results, supported by a theoretical foundation, show that the effective size and success of the collective are highly dependent on properties of the learning algorithm. This highlights the necessity of taking the learning algorithm into account when studying the impact of collective action in machine learning.
@@ 1 Introduction
With the rapid increase in deployed machine learning models, a large number of firms rely on data contributed by users (Gerlitz & Helmond, 2013) to train their algorithms. In response to this, consumers and users have searched for ways to alter their data to influence the outputs of such models (Chen, 2018; Burrell et al., 2019; Rahman, 2021). Algorithmic collective action (Olson, 1965; Hardt et al., 2023) has emerged as a formal framework to study the effect a coordinated group of individuals can have on such models, by altering the data they provide to the firm. This naturally leads to various technical questions regarding how large the collective needs to be, how effective different data altering strategies are, and which details about firm’s algorithm can be leveraged by the collective to achieve the best results.
@@
Hardt et al. (2023) initiated the formal study into the success of different collective action strategies. Their work provided theoretical analysis based on the fractional size of the collective and on the properties of the signal with regards to the original data distribution. However, their results only hold for Bayes (sub-)optimal classifiers which do not immediately adapt to peculiarities of practical learning algorithms. In this work, we investigate two types of commonly used learning algorithms that exhibit distinct properties: (1) Distributional robustness and (2) Simplicity bias. We show how these properties lead to significantly different levels of collective success that are unexplained by prior work.
@@
When the data is composed of multiple sub-populations, algorithms that optimise for average performance often perform poorly in minority sub-populations (Meinshausen & Bühlmann, 2015). A fairness-focused firm will want to ensure that their trained learning model performs well uniformly on all sub-populations, as opposed to being on-average good. Distributionally Robust Optimisation (DRO) is a family of algorithms designed to maximize this per-group accuracy (Hashimoto et al., 2018; Wang et al., 2020). We show that as a consequence, a small collective achieves higher success when the training algorithm performs DRO, compared to standard empirical risk minimization (ERM). Conversely, a large collective in the same settings achieves lower success, contradicting the expectation set by previous work regarding the effectiveness of large collectives.
@@
Second, most machine learning algorithms used today are based on some form of gradient descent (GD). Such algorithms, including the popular Stochastic Gradient Descent (SGD), Adam (Kingma & Ba, 2015), and RMSProp, exhibit a preference for learning functions that are “simpler". This inductive bias of GD algorithms is popularly referred to as simplicity bias (Kalimeris et al., 2019; Shah et al., 2020). In practice, this preference results in the model “overlooking” certain complex features. We demonstrate that these overlooked features can be leveraged by a collective to design a strategy that will gain higher success compared to what is possible on a Bayes optimal classifier.
@@ 2.1 Collective Action
While they are sufficient to characterise the success of the collective for a Bayes (sub-)optimal learner, various practically deployed algorithms show different behaviours, as is discussed in Section 3 and 4.
@@ 2.2 Distributionally Robust Optimisation
In Section 3, we inspect collective action on a set of learning algorithms that target Distributionally-Robust objectives (Delage & Ye, 2010; Sagawa et al., 2020; Duchi & Namkoong, 2021). Intuitively, these algorithms aim to learn classifiers that perform equally well on a set of distributions as opposed to any single one. Formally, they minimise the following objective
@@
where ${\mathcal{Q}}_{p}$ is an uncertainty set of distributions close to $p$ over which we want to control the risk, $\ell$ is the loss function, and $g_{\theta}$ is a function with parameters $\theta$. There are many possible definitions of the uncertainty set that the algorithm is controlling for. One possible choice is an $f$-divergence111 Definition 2 in the appendix. ball (Ali & Silvey, 1966; Csiszár, 1967).
@@
Thus, DRO algorithms are often employed when there is concern about performance on “similar” distributions or on subgroups of the data (Namkoong & Duchi, 2016; Duchi & Namkoong, 2019). In Section 3, we investigate how the success of the collective changes when minimising ${\mathcal{R}}_{\mathrm{dro}}(\theta)$ as opposed to performing simple Empirical Risk Minimisation (ERM).
@@ 5 Conclusion and Future Work
In this work we conducted a theoretical and empirical study on how the choice of learning algorithms affects the success of a collective action to plant a signal. We presented various approaches a collective can take to design their signal when they have knowledge about the learning algorithm. In particular, these approaches include maintaining a small collective size against algorithms like JTT and LfF, influencing the validation set against CVaR-DRO, and changing the complexity of the signal against SGD. While we have focused on these three algorithms, these phenomena are relevant for several popular algorithms. Common algorithms in the topic of domain adaptation (Koh et al., 2021), distribution shift (Hendrycks & Dietterich, 2019), and improving fairness (Berk et al., 2017) use some form of DRO algorithms. It would be interesting to consider the impact of other kinds of algorithms used to improve worst group performance including un-supervsied and self-supervised representation learning (Shi et al., 2023) and fairness-inducing in-processing (Prost et al., 2019) and post-processing algorithms (Ţifrea et al., 2024). In a similar vein, recent research has seen a surge in algorithms that aim to improve safety of outputs of generative models including techniques like adversarial training (Ziegler et al., 2022), DPO (Rafailov et al., 2024), and RLHF (Christiano et al., 2017). It is important to also consider the impact of these algorithms on the success of a well-designed collective.
@@
Algorithmic biases akin to simplicity bias are also exhibited by SGD training of various state-of-the art deep neural networks. This includes a texture bias for CNN (Hermann et al., 2020), an in-context bias for language models (Levine et al., 2022), and word-order biases in LSTMs and transformers (White & Cotterell, 2021). Our work discusses how these biases can affect the design of the collective signal and how successful they are expected to be. However, several algorithms also display a different kind of biases. For example, differentially private algorithms are unable to learn minority sub-populations (Bagdasaryan et al., 2019; Sanyal et al., 2022) and adversarially robust algorithms (Madry et al., 2018) are used to improve the smoothness of learned models. One direction for future work is to consider the impact of these algorithmic biases on the success of the collective.
@@
Finally, a third important direction of future research is understanding what level of information and influence is required by the collective to design their signal. For example, our results suggest that improving success against certain DRO algorithms requires less information about the learning algorithm compared to success against SGD style algorithms. However, for iterative re-weighting algorithms like CVar-DRO (Sagawa et al., 2020), it is important to influence both the validation set and the training set. Future work should investigate whether a uniformly randomly selected collective can be as powerful as a strategically chosen collective (e.g. in adversarial vulnerability (Paleka & Sanyal, 2023)). We hope that our work will inspire further research into practical and algorithm-dependent view on collective action.